OpenStudy (anonymous):

Determine whether the polynomial is a difference of squares and if it is, factor it. y2 – 16 A. is a difference of squares: (y – 4)2 B. is a difference of squares: (y – 4)(y + 4) C. is not a difference of squares D. is a difference of squares: (y – 8)(y + 8)

OpenStudy (a1234):

Well, y^2 and 16 are perfect squares. So it is a difference of square since you have the negative. Keep in mind that 4 * 4 = 16, while -4 * 4 = -16. See your answer now?

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

\[a^{2}-b^{2}=(a-b)(a+b)\] Let: \[a=y\] and \[b=4\] then \[a^{2}-b^{2}=y^{2}-4^{2}=(y-4)(y+4)\] B

OpenStudy (anonymous):

Thanks!